OpenStudy (anonymous):

Electric charge is distributed over the disk x^2 + y^2 ≤ 11 so that the charge density at (x,y) is σ(x,y) = 19 + x^2 + y^2 coulombs per square meter. Find the total charge on the disk. ?????

OpenStudy (turingtest):

the disk is radius\[r=\sqrt11\]converting to polar coordinates\[x=r \cos \theta\]\[y=r \sin \theta\]making these substitutions gives\[\sigma (x,y)=19+11(\cos^2 \theta +\sin^2 \theta)=21\]in polar coordinates\[dA=rdrd \theta\]so our integral is now\[\int\limits_{0}^{2 \pi}\int\limits_{0}^{\sqrt11}21rdrd \theta=\int\limits_{0}^{2 \pi}21{(\sqrt11)^2\over2}d \theta=\int\limits_{0}^{2 \pi}{231\over2}d \theta=231 \pi\]

OpenStudy (anonymous):

so 341pi !?

OpenStudy (turingtest):

Yet another slight typo... 19+11=30 put 30 where I have 21 and it will be correct sorry again...

OpenStudy (anonymous):

i meant 330 !

OpenStudy (anonymous):

yeah but i plug it in and its not correct !

OpenStudy (turingtest):

really? hmm.. let me see

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

I guess you have to keep r as a variable since you integrate across it, so\[σ(x,y)=19+r^2(\cos^2θ+\sin^2θ)=19+r^2\]\[\int\limits_{0}^{2\pi}\int\limits_{0}^{\sqrt11}19r+r^3drd \theta={539\pi \over2}\] Please let me know if this is right.

OpenStudy (turingtest):

finally!

OpenStudy (anonymous):

awesome thanks ! it makes sense too !

OpenStudy (turingtest):

my bad earlier, I apologize.

OpenStudy (anonymous):

haha no man thank u for helping me out ! its math we all do mistakes !